# Symmetric Linear Bandits with Hidden Symmetry

### v1

@@ 4.1 The Explore-Models-then-Commit Algorithm
Algorithm 1 may incur significant computational costs, as it concurrently updates all $M$ models. Nevertheless, this process can be entirely parallelised using tools such as Ray [1]. As a result, the algorithm’s runtime will remain constant even as the size $M$ of $\mathcal{M}$ increases.
@@ 4.2.1 Assumptions
The partition corresponding to $\mathcal{G}$ belongs to a small subclass of partitions $\mathcal{Q}_{d,\leq d_{0}}\subset\mathcal{P}_{d,\leq d_{0}}$. In particular, $\pi_{\mathcal{G}}\in\mathcal{Q}_{d,d_{\star}},\;\text{for some $d_{\star}\leq d_{0}$}$, and for each $k\in[d_{0}]$, $|\mathcal{Q}_{d,k}|\leq O(d^{k})$.
Assumption 6 (Exploratory distribution).
There exists a sub-Gaussian exploratory distribution $\nu\in\Delta(\mathcal{X})$. Specifically, suppose there is a random vector $Z$ whose distribution is $\nu$, then for any $a\in\mathbb{S}^{d-1}$, $Z,a\right>$ is a sub-Gaussian random variable.
Assumption 7 (Boundedness).
For all $x\in\mathcal{X}$, $\left\|x\right\|_{2}\leq 1$ and |$x,\theta_{\star}\right>\right|\leq R_{\mathrm{\max}}$, for some $R_{\mathrm{\max}}>0$.
As a consequence of Assumption 5, the cardinality of the collection of fixed point subspaces is not too large, particularly, $M=O(d^{d_{0}})$. This assumption holds for broad classes of pattern-avoiding partitions [29], such as interval partitions and non-crossing partitions. We refer the readers to Appendix D for a detailed discussion.
@@
A set partition of $[d]$ is an interval partition or partition of interval if its parts are interval. We denote $\mathcal{I}_{d}$ as the collection of all interval partition of $d$. $\mathcal{I}_{d}$ admits a Boolean lattice of order $2^{d-1}$, making it equivalent to the sparsity structure in $d-1$ dimensions. Specifically, consider the set of entries of parameters $\varphi\in\mathbb{R}^{d}$ with a linear order, that is, $\varphi_{1}<\varphi_{2}<\dots<\varphi_{d}$. Then define the variable $\theta\in\mathbb{R}^{d-1}$ such that $\theta_{i}=(\varphi_{i+1}-\varphi_{i})$. Each interval partition on the entries of $\varphi$ will determine a unique sparse pattern of $\theta$. Therefore, it is clear that the cardinality of the set of interval partition with $d_{0}$ classes is bounded as $|\mathcal{I}_{d,\leq d_{0}}|=O(d^{d_{0}})$. Moreover, as a result, symmetric linear bandit is strictly harder than sparse bandit and inherits all the computational complexity challenges of sparse linear bandit, including the NP-hardness of computational complexity.
@@
Proposition 9 (Regret lower bound [21]).
Consider symmetric linear bandits in which Assumption 5, 6, 7 hold. Then any bandit algorithm must suffer regret $\mathbf{R}_{T}=\Omega\left(\min\left(d_{0}^{\frac{1}{3}}T^{\frac{2}{3}},\sqrt{d_{0}dT}\right)\right)$.
@@ 4.2.2 Main Regret Upper Bound Result
Theorem 10.
Suppose the Assumptions 5, 6, 7 hold. With the choice of $t_{1}=\Omega(d_{0}\log(dT))$, then the regret of Algorithm 1 is upper bounded as
$\mathbf{R}_{T}=O\left(R_{\mathrm{\max}}^{\frac{1}{3}}T^{\frac{2}{3}}\sigma^{\frac{2}{3}}(d_{0}\log(dT))^{\frac{1}{3}}\right).$ (6)
@@
Suppose the Assumptions 5, 6, 7 hold. For $t_{1}=\Omega(d_{0}\log(d/\delta))$, with probability at least $1-\delta$, one has the estimate
$\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}=O\left(\sqrt{\frac{\sigma^{2}d_{0}\log(d/\delta)}{t_{1}}}\right).$ (7)
The constants hidden in asymptotic notations are independent of $d,\;d_{0},\;T$.
@@
We provide a proof sketch here and defer their full proof to Appendix B. Our proof borrows techniques from Gaussian model selection [19] and the compressed sensing literature [8]. There are two steps to bound the risk error as in Lemma 11:
@@
$\left\|\widehat{\bm{f}}_{\widehat{m}}-\bm{f}_{\star}\right\|_{2}^{2}\leq C\sigma^{2}\log\left(\frac{M}{\delta}\right).$ (8)
@@
To bound the risk error from the prediction error, we invoke the restricted isometry property on the union of subspaces of a sub-Gaussian random matrix as in [8]. For more detail, let us define the extension of the collection $\mathcal{M}$ as
¯:=$\,\overline{\!{\mathcal{M}}}:=\left\{\mathrm{conv}\left(m\cup m^{\prime}\right)\mid\;m,\;m^{\prime}\in\mathcal{M}\right\},$ (9)
where $\mathrm{conv}(S)$ is the convex hull of the set $S\subset\mathbb{R}^{n}$. We have that $\,\overline{\!{\mathcal{M}}}$ is a collection of subspaces, that is, $\mathrm{conv}\left(m\cup m^{\prime}\right)$ is indeed a subspace [8]. Denote $\,\overline{\!{M}}:=|\,\overline{\!{\mathcal{M}}}|$, then we have $\,\overline{\!{M}}=(M^{2}-M)/2$. Moreover, if dimension of subspace in $\mathcal{M}$ is at most $d_{0}$, then the dimension of subspace in $\,\overline{\!{\mathcal{M}}}$ is at most $2d_{0}$. The reason of introducing the extension $\,\overline{\!{\mathcal{M}}}$ is that $\widehat{\theta}$ and $\theta_{\star}$ can belong to two different subspaces of $\mathcal{M}$, and $\widehat{\theta}-\theta_{\star}$ may not lie in any subspace of $\mathcal{M}$, but in $\,\overline{\!{\mathcal{M}}}$.
An important property of the design matrix $X$, which allows one to recover $\theta_{\star}$ with the knowledge that $\theta_{\star}$ is in a subspace $m\in\mathcal{M}$, can be captured by the following notion of restricted isometry property (RIP):
@@
For any matrix $X$ and any collection of subspaces $\,\overline{\!{\mathcal{M}}}$, we define $\,\overline{\!{\mathcal{M}}}$-restricted isometry constant $\delta_{\,\overline{\!{\mathcal{M}}}}(X)$ to be the smallest quantity such that
$(1-\delta_{\,\overline{\!{\mathcal{M}}}}(X))\leq\frac{\left\|X\theta\right\|_{2}^{2}}{\left\|\theta\right\|_{2}^{2}}\leq(1+\delta_{\,\overline{\!{\mathcal{M}}}}(X)).$ (10)
We have the minimum number of samples required so that a random matrix $X$ satisfies RIP for a given constant $\delta_{\,\overline{\!{\mathcal{M}}}}(X)$ with high probability as follows.
Proposition 15 ([8]’s Corollary 3.4).
Suppose $X=[x_{t}]_{t\in[n]}$, where $x_{t}$’s are is i.i.d. sub-gaussian random vectors, and let
$n\geq\frac{2}{c\delta_{\,\overline{\!{\mathcal{M}}}}(X)}\left(\log(2\,\overline{\!{M}})+d_{0}\log\left(\frac{12}{\delta_{\,\overline{\!{\mathcal{M}}}}(X)}\right)+\log(\delta^{-1})\right),$ (11)
for some constant $c>0$ only depends on the distribution of $x_{t}$. Then, with probability at least $1-\delta$, and for any $\theta_{1},\theta_{2}$ in subspaces of $\mathcal{M}$, one has that
$(1-\delta_{\,\overline{\!{\mathcal{M}}}}(X))\left\|\theta_{1}-\theta_{2}\right\|_{2}^{2}\leq\frac{1}{n}\left\|X(\theta_{1}-\theta_{2})\right\|_{2}^{2}\leq(1+\delta_{\,\overline{\!{\mathcal{M}}}}(X))\left\|\theta_{1}-\theta_{2}\right\|_{2}^{2}$ (12)
Moreover, if $(x_{t})_{t\in[n]}$ are i.i.d. normal random vectors, then $c=7/18$, see [8]. Lemma 11 is followed by combining Proposition 15 and Proposition 13.
5 Improved Regret Upper Bound
We now show that by adding more structure (i.e., well-separatedness) to the setting, we can further improve the regret upper bound to $O(\sqrt{T})$ (Section 5.1). We then discuss an interesting open problem in Section 5.2.
5.1 Symmetric Linear Bandits with Well-Separated Partitions
[…]
@@ 5.1 Symmetric Linear Bandits with Well-Separated Partitions
$\left\|\theta_{\star}-\Pi_{m}(\theta_{\star})\right\|_{2}^{2}\geq\frac{\varepsilon_{0}^{2}}{2}.$ (13)
@@
Lemma 17.
Suppose the Assumptions 5, 6, 7, 16 hold. After $t_{2}=\Omega\left(\frac{\sigma^{2}d_{0}\log(d\delta^{-1})}{\varepsilon_{0}^{2}}\right)$ exploring rounds, the Algorithm 2 returns $\widehat{m}\ni\theta_{\star}$, with probability at least $1-\delta$.
Combining Lemma 17 with the standard regret bound $\tilde{O}(d_{0}\sqrt{T})$ for linear bandit algorithms [2] on the chosen subspace $\hat{m}$, we obtain the following regret bound:
@@
Suppose the Assumptions 5, 6, 7, 16 hold. Let $t_{2}=\Omega\left(\frac{\sigma^{2}d_{0}\log(dT)}{\varepsilon_{0}^{2}}\right)$. Then the regret of Algorithm 2 is upper bounded as
$\mathbf{R}_{T}=O\left(\frac{R_{\mathrm{\max}}\sigma^{2}d_{0}\log(dT)}{\varepsilon_{0}^{2}}+\sigma d_{0}\sqrt{T\log(d_{0}T)}\right).$ (14)
That is, if the separating constant $\varepsilon_{0}$ is known in advance and $\varepsilon_{0}\geq T^{-1/4}$, then we can achieve $O(d_{0}\sqrt{T}\log(dT))$ regret upper bound.
5.2 Adapting to the Separating Constant $\varepsilon_{0}$
A weakness of Algorithm 2 is that without knowing that $\varepsilon_{0}\geq T^{-1/4}$ is true a priori, there may be possible mis-specification error, which leads to linear regret if one applies the algorithm naively. On the other hand, Algorithm 1 can always achieve regret $O(T^{2/3})$ in the worst case. As such, the following question arises: Does there exist an algorithm that, without the knowledge of $\varepsilon_{0}$, can achieve regret $O(\sqrt{T})$ whenever $\varepsilon_{0}\geq T^{-1/4}$, but guarantees the worst-case regret of $O(T^{2/3})$?
We now propose a simple method which has $O(\sqrt{T})$ regret whenever the separating constant is large, and enjoys a worst-case regret guarantee of $O(T^{3/4})$ (slightly worse than $O(T^{2/3})$). We refer the reader to Appendix C.2 for a detailed description of the algorithm and further discussion.
Corollary 19.
Suppose the Assumptions 5, 6, 7 hold. Then, there exists an algorithm which achieves regret bound as follows:
  • (i)
  • (ii)
    [None well-separated partitions] If $\varepsilon_{0}<T^{-1/4}$, then $\mathbf{R}_{T}=\tilde{O}(d_{0}\sqrt{T}+T^{\frac{3}{4}}\sqrt{d_{0}})$.

### v2

@@ 4.1 The Explore-Models-then-Commit Algorithm
The key step of Algorithm 1 that may incur significant costs is solving equation (4) (line 6). Without additional information about $\mathcal{M}$, one might need to enumerate all models in $\mathcal{M}$ and optimize among them, which would induce a time complexity of $O(nd^{cd_{0}})$. However, if we have more information about the partitions, e.g., if they are non-crossing or non-nesting partitions, their lattice structures can be exploited to speed up the optimization process of solving equation (4). Due to space limitations, we refer readers to Appendix D.3 for a detailed explanation of a subroutine that leverages these lattice structures for more efficient computation. Additionally, Section 6 demonstrates that our Algorithm 1, when using the lattice search algorithm for non-crossing partitions and non-nesting partitions as a subroutine, achieves polynomial computational complexity of $O(nd^{5})$ and guarantees low regret.
@@ 4.2.1 Assumptions
The partition corresponding to $\mathcal{G}$ belongs to a small subclass of partitions $\mathcal{Q}_{d,\leq d_{0}}\subset\mathcal{P}_{d,\leq d_{0}}$. In particular, $\pi_{\mathcal{G}}\in\mathcal{Q}_{d,d_{\star}},\;\text{for some $d_{\star}\leq d_{0}$}$, and for each $k\in[d_{0}]$, there exists a constant $c>0$, such that $|\mathcal{Q}_{d,k}|\leq O(d^{ck})$.
Assumption 6 (Bounded set of arms).
There are positive numbers $K_{x},R_{\mathrm{\max}}$, such that, for all $x\in\mathcal{X}$ and $m\in\,\overline{\!{\mathcal{M}}}$, $\left\|\Pi_{m}(x)\right\|_{2}^{2}\leq K_{x}$, and |$x,\theta_{\star}\right>|\leq R_{\mathrm{\max}}$.
As a consequence of Assumption 5, the cardinality of the collection of fixed point subspaces is not too large, particularly, $M=O(d^{cd_{0}})$. First, we note that this class includes interval partitions, a structure equivalent to sparsity as a strict subset, as explained below.
@@
A set partition of $[d]$ is an interval partition or partition of interval if its parts are interval. We denote $\mathcal{I}_{d}$ as the collection of all interval partition of $d$. $\mathcal{I}_{d}$ admits a Boolean lattice of order $2^{d-1}$, making it equivalent to the sparsity structure in $d-1$ dimensions. Specifically, consider the set of entries of parameters $\varphi\in\mathbb{R}^{d}$ with a linear order, that is, $\varphi_{1}\geq\varphi_{2}\geq\dots\geq\varphi_{d}$. Then define the variable $\theta\in\mathbb{R}^{d-1}$ such that $\theta_{i}=(\varphi_{i}-\varphi_{i+1})$. Each interval partition on the entries of $\varphi$ will determine a unique sparse pattern of $\theta$. Therefore, it is clear that the cardinality of the set of interval partition with $d_{0}$ classes is bounded as $|\mathcal{I}_{d,\leq d_{0}}|=O(d^{d_{0}})$. Moreover, as a result, symmetric linear bandit is strictly harder than sparse bandit and inherits all the computational complexity challenges of sparse linear bandit, including the NP-hardness of computational complexity.
Figure 1: Partition that respects the underlying ordered tree.
Apart from sparsity, class of partitions with sub-exponential size also naturally appears when there is a hierarchical structure on the set $[d]$, and the partitioning needs to respect this hierarchical structure. A partition that respects an ordered tree groups the children of the same node into a single equivalence class, for example, see Figure 1. It is shown in [15] and the cardinality of the set of partitions that respect ordered trees is sub-exponential. Furthermore, as shown in [15] there is a bijection between partitions that respect ordered trees and the set of non-crossing partitions.
A real-life example that meets these assumptions is the subcontractor example in the introduction: A hierarchical structure may exist, where a hired subcontractor can further subcontract parts of the work to others. A tree represents the hierarchical order among subcontractors, where subcontractors hired by another contractor can be grouped into one class. Further real-life examples of non-crossing partitions and other structured partitions that satisfy sub-exponential cardinality, such as non-nesting partitions and pattern-avoidance partitions [33], can be found in Appendix D.2.
Next, similar as [23], we define the exploratory distribution as follows.
Definition 8 (Exploratory distribution).
The exploratory distribution $\nu\in\Delta(\mathcal{X})$ is solution of the following optimisation problem
$\nu=\operatorname*{arg\,max}_{\omega\in\Delta(\mathcal{X})}\lambda_{\min}\left(\mathbb{E}_{x\sim\omega}[xx^{\top}]\right),\quad V:=\mathbb{E}_{x\sim\nu}[xx^{\top}],\quad C_{\min}(\mathcal{X}):=\lambda_{\min}(V).$ (5)
@@
Proposition 9 (Regret lower bound).
There exist symmetric linear bandit instances in which Assumption 5, 6 hold with $K_{x}=8d_{0}$, such that, any bandit algorithm must suffer regret $\mathbf{R}_{T}=\Omega\left(\min\left(C_{\min}(\mathcal{X})^{-\frac{1}{3}}d_{0}^{\frac{2}{3}}T^{\frac{2}{3}},\sqrt{dT}\right)\right)$.
@@ 4.2.2 Main Regret Upper Bound Result
Theorem 10 (Regret upper bound).
Suppose the Assumptions 5, 6 hold. With the choice of $t_{1}=R_{\mathrm{\max}}^{-\frac{2}{3}}\sigma^{\frac{2}{3}}C_{\min}^{-\frac{1}{3}}(\mathcal{X})K_{x}^{\frac{1}{3}}d_{0}^{\frac{1}{3}}T^{\frac{2}{3}}(\log(dT))^{\frac{1}{3}}$, then the regret of Algorithm 1 is upper bounded as
$\mathbf{R}_{T}=O\left(R_{\mathrm{\max}}^{\frac{1}{3}}\sigma^{\frac{2}{3}}C_{\min}^{-\frac{1}{3}}(\mathcal{X})K_{x}^{\frac{1}{3}}d_{0}^{\frac{1}{3}}T^{\frac{2}{3}}(\log(dT))^{\frac{1}{3}}\right)$ (6)
Remark 11.
We note that when $K_{x}=O(d_{0})$, as in the lower bound instances, our upper bound is $\tilde{O}\left(C_{\min}^{-\frac{1}{3}}(\mathcal{X})d_{0}^{\frac{2}{3}}T^{\frac{2}{3}}\right)$, which matches the lower bound in Proposition 9.
@@
Suppose the Assumptions 5, 6 hold. For $t_{1}=\Omega(K_{x}^{2}d_{0}C_{\min}^{-2}(\mathcal{X})\log(d/\delta))$, with probability at least $1-\delta$, one has the estimate
$\left\|\theta_{\star}-\widehat{\theta}_{t_{1}}\right\|_{2}=O\left(\sqrt{\frac{\sigma^{2}d_{0}\log(d/\delta)}{C_{\min}(\mathcal{X})t_{1}}}\right).$ (7)
@@
We provide a proof sketch here and defer their full proof to Appendix B.1. Our proof borrows techniques from Gaussian model selection [21] and the compressed sensing literature [8]. There are two steps to bound the risk error as in Lemma 12:
@@
$\left\|\widehat{\bm{f}}_{\widehat{m}}-\bm{f}_{\star}\right\|_{2}^{2}\leq C\sigma^{2}\log\left(M\delta^{-1}\right).$ (8)
@@
To bound the risk error from the prediction error, we invoke the restricted isometry property on the union of subspaces of a sub-Gaussian random matrix as in [8]. Note that $\widehat{\theta}$ and $\theta_{\star}$ can belong to two different subspaces of $\mathcal{M}$, and $\widehat{\theta}-\theta_{\star}$ may not lie in any subspace of $\mathcal{M}$, but in $\,\overline{\!{\mathcal{M}}}$. An important property of the design matrix $X$, which allows one to recover $\theta_{\star}$ with the knowledge that $\theta_{\star}$ is in a subspace $m\in\mathcal{M}$, can be captured by the following notion of restricted isometry property (RIP):
@@
For any matrix $X$, any collection of subspaces $\,\overline{\!{\mathcal{M}}}$ and any $\theta\in m\in\,\overline{\!{\mathcal{M}}}$, we define $\,\overline{\!{\mathcal{M}}}$-restricted isometry constant $\delta_{\,\overline{\!{\mathcal{M}}}}(X)$ to be the smallest quantity such that
$(1-\delta_{\,\overline{\!{\mathcal{M}}}}(X)){\left\|\theta\right\|_{2}^{2}}\leq\left\|X\theta\right\|_{2}^{2}\leq(1+\delta_{\,\overline{\!{\mathcal{M}}}}(X)){\left\|\theta\right\|_{2}^{2}}.$ (9)
We have the minimum number of samples required so that a random matrix $X$ satisfies RIP for a given constant with high probability as Proposition 16 below. Then, Lemma 12 is followed by combining Proposition 16 and Proposition 14.
Proposition 16.
Let $X=[x_{t}]_{t\in[n]}$, where $x_{t}$’s are is i.i.d. drawn from $\nu$, and let $n=\Omega\left(C_{\min}^{-2}(\mathcal{X})K_{x}^{2}\left(\log(2\,\overline{\!{M}}d_{0}\delta^{-1})\right)\right).$ Then, with probability at least $1-\delta$, and for any $\theta_{1},\theta_{2}$ in subspaces of $\mathcal{M}$, one has that
$\left\|\theta_{1}-\theta_{2}\right\|_{2}^{2}\leq 2C_{\min}^{-1}(\mathcal{X})n^{-1}\left\|X(\theta_{1}-\theta_{2})\right\|_{2}^{2}.$ (10)
5 Improved Regret Upper Bound with Well-Separated Partitions
We now show that by adding more structure (i.e., well-separatedness) to the setting, we can further improve the regret upper bound to $O(\sqrt{T})$. In particular, we will introduce the notion of well-separatedness in this section, and show that this notion can lead to improved (i.e., $O(\sqrt{T})$) regret bounds.
@@ 5 Improved Regret Upper Bound with Well-Separated Partitions
$\left\|\theta_{\star}-\Pi_{m}(\theta_{\star})\right\|_{2}^{2}\geq\varepsilon_{0}^{2}/2.$ (11)
@@
Suppose the Assumptions 5, 6, 17 hold. Let $t_{2}=\Omega\left(\frac{\sigma^{2}K_{x}^{2}d_{0}\log(dT)}{C_{\min}^{2}(\mathcal{X})\varepsilon_{0}^{2}}\right)$. Then, Algorithm 2 returns $\widehat{m}\ni\theta_{\star}$ with probability at least $1-1/T$, and its regret is upper bounded as
$\mathbf{R}_{T}=O\left(\frac{R_{\mathrm{\max}}\sigma^{2}K_{x}^{2}d_{0}\log(dT)}{C_{\min}^{2}(\mathcal{X})\varepsilon_{0}^{2}}+\sigma d_{0}\sqrt{T\log(K_{x}T)}\right).$ (12)
That is, if the separating constant $\varepsilon_{0}$ is known in advance and $\varepsilon_{0}\geq T^{-1/4}$, then we can achieve $O(d_{0}\sqrt{T}\log(K_{x}T))$ regret upper bound.
Remark 19.
A weakness of Algorithm 2 is that without knowing that $\varepsilon_{0}\geq T^{-1/4}$ is true a priori, there may be possible mis-specification error, which leads to linear regret if one applies the algorithm naively. On the other hand, Algorithm 1 can always achieve regret $O(T^{2/3})$ in the worst case. As such, the following question arises: Does there exist an algorithm that, without the knowledge of $\varepsilon_{0}$, can achieve regret $O(\sqrt{T})$ whenever $\varepsilon_{0}\geq T^{-1/4}$, but guarantees the worst-case regret of $O(T^{2/3})$? Toward answering this question, we propose a simple method which has $O(\sqrt{T})$ regret whenever the separating constant is large, and enjoys a worst-case regret guarantee of $O(T^{3/4})$ (slightly worse than $O(T^{2/3})$). We refer the reader to Appendix C.2 for a detailed description of the algorithm, its regret bound and further discussion.
6 Experiment
To illustrate the performance of our algorithm, we conduct simulations where the entries of $\theta_{\star}$ satisfy three cases: sparsity, non-crossing partitions and non-nesting partitions. We refer readers to Appendix D.1 for a more formal description of non-crossing partitions, non-nesting partitions, and why the interval partition (i.e., the partition structure equivalent to sparsity) is a strict subset of both non-crossing and non-nesting partitions. Since sparsity is equivalent to a strict subset of non-crossing and non-nesting partitions, we compare our Algorithm 1 with the sparse-bandit ESTC algorithm proposed in [23] as a benchmark in all environments. The set of arms $\mathcal{X}$ is $\sqrt{d}\mathbb{S}^{d-1}$, $\sigma=0.1$, and $d=100$, $d_{0}=15$. The ground-truth sparse patterns, partitions and $\theta_{\star}$ are randomized before each simulation.
The regret of both algorithms is shown in Figure 2, which indicates that our algorithm performs competitively in the sparsity case and significantly outperforms the sparse-bandit algorithm in cases of non-crossing and non-nesting partitions. Due to space limitations, we refer the reader to Appendix E for a detailed description of the experiments, including how we applied the sparse-bandit algorithm in the cases of non-crossing and non-nesting partitions, and how we ran Algorithm 1 in the case of sparsity. Additionally, we explain how we exploited the particular structure of non-crossing and non-nesting partitions to enable efficient computation in Appendix E.
Figure 2: Regret of EMC (Algorithm 1) and of ESTC proposed in [23], in cases of sparsity, non-crossing partitions, and non-nesting partitions.